A population of students were polled: Do you like chocolate donuts or glazed? 20% said neither, 70% liked chocolate and 60% liked glazed. What percent liked both chocolate and glazed?

Answer by venugopalramana(3286)

A population of students were polled: Do you like chocolate donuts or glazed? 20% said neither, 70% liked chocolate and 60% liked glazed. What percent liked both chocolate and glazed?
LET TOTAL BE = 100
LIKED NEITHER = 20
HENCE THOSE WHO LIKED =80
LIKED CHACOLTE=70
LIKED GLAZED =60
TOTA NUMBER =70+60 = 130
BUT THOSE WHO LIKED ARE ONLY = 80
HENCE THOSE WHO LIKED BOTH OR THE OVER LAP = 130-80 = 50...THAT IS 50%